# Large Language Model-Enabled Multi-Agent Manufacturing Systems

### v1

@@ I Introduction
The escalating demand for customized products and the increasing complexities of production processes are driving a transformation in the manufacturing sector. This paradigm shift has required manufacturers to quickly adjust to evolving product specifications and adeptly address unforeseen operational changes. In this context, manufacturers are required to be not only agile and responsive but also to maintain high-quality products and cost efficiency in a dynamically changing environment.
@@
One significant approach for enhancing the adaptability and quick response of manufacturing systems is the use of multi-agent control [1, 2]. A Multi-Agent System (MAS) is composed of several autonomous agents for decision-making, communication, and coordination [3]. Many multi-agent architectures have been developed for controlling manufacturing systems [1, 4, 5, 6]. By enabling decentralized decision-making and enhancing coordination among various agents, MAS have contributed significantly to handling complexities in the manufacturing system [7].
@@
However, there are still challenges in integrating MAS frameworks in the manufacturing context. While agents in MAS can automate processes and make decisions from available data, their ability to adapt to new formats or specifications is limited, leading to increased downtime. Thus, a significant challenge includes the limited ability of agents to dynamically adapt to new information including product specifications (RCh1). Another challenge lies in processing unstructured textual data. Textual data is crucial because it is more expressive than numerical data, offering vast knowledge that is currently underutilized in manufacturing, where structured data fails to include varied influencing factors [8]. The capability of natural language processing to formalize unstructured textual data can enable downtime reduction and process optimization [8]. Therefore, advanced natural language processing is essential for interpreting and implementing textual data from human operators or interconnected systems (RCh2).
@@
The advent of Large Language Models (LLMs) presents a novel solution to these challenges. LLMs, such as OpenAI’s GPT-3 [9] and GPT-4 [10], provides opportunities to bring more flexibility and adaptability into multi-agent manufacturing systems. With Machine to Machine (M2M) protocols in the manufacturing industry, such as Message Queuing Telemetry Transport (MQTT) [11] and MTConnect [12] that uses predefined structures, incorporating LLMs offer benefits. LLMs can interpret, generate, and act on natural language instructions, facilitating complex decision-making processes and adaptive responses to changing manufacturing conditions [13]. The integration of natural language capabilities may also benefit operators, engineers, and other system users who have limited experience with the specific process or system [14, 15]. For manufacturing systems, natural language capabilities can allow new users to quickly and efficiently adapt to managing and optimizing processes and resources [16]. The main contributions of this paper include: Leveraging LLM to allocate and execute manufacturing tasks based on user instructions, to enhance task efficiency (Cn1), enabling LLM agents to interpret and adapt to evolving specifications ensuring the production process is correctly executed with the appropriate resources (Cn2), and applying advanced natural language processing capabilities to enable agents to effectively comprehend context-specific instructions from users and other agents (Cn3).
@@
The rest of the manuscript is organized as follows. Section II provides background regarding MAS, LLMs in MAS, and LLMs in manufacturing. Section III details the architecture and functioning of the proposed framework. Section IV demonstrates a case study to illustrate the applications and effectiveness of our framework. Section V presents challenges and limitations encountered while integrating LLM. Section VI summarizes the paper and discusses future work.
@@ II-A Multi-Agent Systems in Manufacturing
Various MAS architectures have been developed to enhance agent communication, significantly contributing to the increased flexibility of manufacturing systems. For instance, the Axiomatic Design of a Multi-Agent Reconfigurable Mechatronic System (ADMARMS) framework is an early MAS architecture designed for reconfigurable manufacturing, incorporating axiomatic design and quantitative methodologies to facilitate design-based reconfigurability. Studies by Wang et al. [17] and Kim et al. [18] have expanded the scope of MAS by integrating these systems with big data analytics and reinforcement learning. Wang et al. focused on improving manufacturing flexibility through agent negotiation for deadlock prevention, while Kim et al. aimed at facilitating autonomous decision-making, particularly for personalized production systems. Additionally, the works of Kovalenko et al. [19] and Bi et al. [6] have concentrated on dynamic resource and task allocation within MAS. Kovalenko et al. introduced a framework for dynamic resource task negotiation, utilizing a network-based coordination strategy that enables Product Agents (PAs) to effectively explore and understand resource capabilities. Similarly, Bi et al. developed a model-based Resource Agent (RA) architecture to address disruptions in manufacturing systems by improving coordination and decision-making capabilities.
While these MAS architectures were developed to enhance agent communication for increased flexibility, they focus on utilizing predefined mathematical models and algorithms for task allocation and agent communication. Furthermore, there has been limited exploration in using natural language interpretation from human operators or other agents to enhance the responsiveness of manufacturing process control. By leveraging LLMs, agents can achieve higher operational flexibility and efficiency in addressing these gaps.
@@ II-B Large Language Models in Multi-Agent Systems
The emergence of LLMs has sparked significant interest in the development of multiple interacting autonomous agents powered by the LLM technology [20]. The agents enabled by LLMs demonstrate enhanced problem-solving capabilities through multi-agent discussions and diverse applications, from sociological studies to complex task solutions. For instance, Park et al. [21] showcase a simulation with 25 generative agents in a virtual town, designed to exhibit realistic human behavior and facilitate studying social understanding through memory retrieval. Another work by Wu et al. [22] introduces an open-source framework called AutoGen that facilitates customizable agent interactions and conversation programming. ChatEval [23] incorporates debater agents in an LLM-powered MAS for text evaluation, achieving enhanced accuracy and depth over traditional methods and single agent setups. LLMs have also been effectively integrated into specialized domains like software engineering. MetaGPT [24], a meta-programming framework, enhances LLM-based multi-agent collaboration using an assembly line approach to assign different roles to each agent for software engineering tasks, outperforming previous chat-based systems in complex problems. Similarly, ChatDev  [25], a virtual chat-powered framework that applies LLMs to automate and streamline the entire software development process, including efficient collaboration across designing, coding, testing, and documenting phases.
These works demonstrate advanced multi-agent interactions, realistic behavior simulation, and efficient tasks, which can be applied in manufacturing to enhance adaptability, coordination, and advanced natural language processing capabilities. Building on the insights from prior studies, we delve into the integration of LLMs in a multi-agent framework, aiming to address the unique requirements and benefits in the manufacturing sector.
@@ II-C Large Language Models in Manufacturing
With the advancement of LLM technology, recent studies have also been increasingly focusing on exploring both the opportunities and challenges associated with implementing LLMs within the manufacturing domain. Makatura et al. [13] explores the application of LLMs in design and manufacturing, including CAD/CAM. Their approach investigates the potential role of LLMs in a variety of tasks, including design specification and manufacturing instruction, and highlights the capabilities and limitations. The analysis of GPT-4 shows strengths in knowledge and iteration in design and manufacturing but faces challenges in reasoning and accuracy, with potential solutions like specialized languages and improved context management [13]. Furthermore, recent research on LLM in manufacturing delves into language models for specific manufacturing applications. Jignasu et al. [26] presents a comprehensive evaluation of LLMs for understanding, debugging, and manipulating G-code in 3D printing, testing their capabilities in error detection and correction, geometric transformations, and discussing their potential and limitations in G-code comprehension. Similar work has been done by Badini et al. [27], which explores using ChatGPT to optimize the G-code generation in additive manufacturing, enhancing efficiency and quality in manufacturing processes. These studies prove that LLM can interpret G-code, essential for tasks like debugging, manipulating, and comprehending the code used in manufacturing processes. Studies by Fakih et al. [28] and Koziolek et al. [29] explore the use of LLMs for enhancing programming in industrial control systems, specifically focusing on Programmable Logic Controllers (PLCs). Fakih et al. developed LLM4PLC, a framework that iteratively refines LLM-generated code through user feedback and verification tools to improve code generation quality. Koziolek et al. conducted an exploratory study to assess GPT-4’s capability to generate control logic from natural language prompts.
Several studies have investigated the integration of LLMs into manufacturing, including areas such as CAD/CAM, G-code, and PLCs, aiming to enhance efficiency in manufacturing processes. However, these studies have not focused on using LLMs with MAS for autonomous task allocation, adaptability to evolving specifications, and interpretation of context-specific instructions. The framework described in Section III focuses on developing autonomous agents that use instructions and specifications in context-specific natural language to dynamically adjust to new product information and coordinate during runtime to complete manufacturing tasks. Our proposed framework can be identified as an initial step toward enhancing the CAD/CAM workflow for automatically adapting CAM processes based on changes in CAD specifications, ensuring seamless manufacturing operations.
@@ III Framework
The novel architecture presented in Figure 1 is designed based on the architecture of a MAS that includes both PAs and RAs. A PA is an agent responsible for decision-making related to a product, while an RA serves as a controller for a specific resource, such as a robot or machine [30].
Figure 1 provides a general overview of an LLM-enabled multi-agent manufacturing system framework, where a user initializes LLM agents with instructions and capabilities. PAs are assigned specific processes including G-code specifications, while possessing the ability to communicate and access historical data for informed decision-making. RAs are responsible for the process execution of tasks, informed by instructions and tool specifications provided by the user. LLM is applied for processing and responding to natural language inputs from both PAs and RAs. For evaluation, the chosen LLM is GPT-4, the latest and most advanced version of GPT developed by OpenAI [10].
@@ III-A Agent Initialization
The agent initialization process aims to configure a cooperative MAS for operation by defining roles, capabilities, and communication protocols for each agent. This setup ensures agents operate effectively and collaborate towards system goals. The following initial prompt is given to all agents in the system:
@@
These instructions ensure that each agent operates within the boundaries of its capabilities and knowledge, engages in effective communication, and contributes to the overall objectives of the MAS.
Following the initial prompts common to all agents in the system, the initialization process involves various parameters, assigning each agent to its specific role. These parameters were selected to achieve the desired behaviors of adaptability, real-time response, autonomous decision-making, and efficient coordination within the MAS. The process includes the following key parameters essential for the agents:
  • Functions: Capabilities within the manufacturing process, that range from retrieving manufacturing history to executing manufacturing tasks.
  • Annotation: Agent’s role and responsibilities, aiding the selection of the appropriate agent among other LLM agents.
@@
  • Inbox: The agent is initialized with an inbox to receive and manage instructions, updates, and feedback from various agents.
@@
The integration of these parameters into LLM-enabled MAS for manufacturing extends the structured and standardized approaches (e.g. FIPA [31] and JADE [32]) to incorporate natural language processing capabilities for more flexible and human-like agent interactions. Integrating user-defined instructions and parameters in agent initialization can be further expanded through user studies to ensure applicability to real scenarios and industries.
Based on the initialization information, the Agent Creation function, cp. Algorithm 1, enables the dynamic generation of agents. The Agent Creation function provides advanced LLM agents with different capabilities in manufacturing systems, improving operational efficiency and adaptability to changing requirements. The algorithm iterates over each entry in the Agent_list, which contains initialization information. For every entry, configuration data configuration for the agent is loaded from a JSON, containing details such as the agent’s name, role, capabilities, and any specific functions the agent can perform. For each agent specified in the configuration data, an empty list function_list is initially prepared. If the ‘functions’ key is present in the agent’s data, the algorithm iterates through each function name listed under this key. For each function name, a reference function_reference is obtained by searching within a provided set of agent functions Agent_functions. If the function is found, its reference is appended to function_list. A new agent instance agent is created with the specified name, agent data, and the list of function references. This agent is then added to the Agents list. This algorithm facilitates the flexible creation of agents allowing for dynamic adaptation to various tasks and scenarios.
@@ III-B Agent Capabilities and LLM Integration
Through the integration of LLM, agents gain the capability to interpret complex instructions and specifications, facilitating a flexible approach to manufacturing operations. The framework employs LLM to enable agents to decode user instructions for manufacturing, integrating a communication mechanism through function information and execution capabilities with the OpenAI Application Programming Interface (API) feature called function calling [33]. This allows for the automatic execution of user-defined functions, including specific manufacturing operations (e.g., milling, drilling) through descriptions that include operation parameters and required inputs like product name, machine name, process name, and product specifications (e.g. G-code).
The ‘chat’ function, cp. Algorithm 2, facilitates dynamic interaction among agents, based on available function information and executing function calls as defined by the manufacturing process requirements, utilizing the appropriate LLM model. The outcome, function_response, includes the details like operation type, product, and machine names, along with a message, msgs, to facilitate the subsequent steps in the manufacturing process. This approach enables context-aware execution of manufacturing tasks, enhancing the adaptability and efficiency of the manufacturing process.
@@ III-C Operational Workflow
The workflow, as depicted in Figure 2, shows how user instructions in textual data can be converted into actionable manufacturing operations. This approach enables the translation of high-level directives of users into machine-executable actions, enhancing the responsiveness of the manufacturing workflow to changing conditions.
The workflow begins with the user initializing PAs and RAs within the manufacturing system. Users describe product goals, sequences of operations, and specifications. This information is then processed by the LLM, which generates responses to guide the process. The PA determines the suitable RA for the next task, prompting it to carry out the required operation. The selected RA constructs the necessary arguments to request the corresponding function call, leading to the execution of the specified process. Upon successful operation, the LLM formulates an appropriate response, informing the PA of the process status. This feedback updates the LLM-integrated PA with the current process status, determining the completion of the task. If additional tasks remain, the cycle repeats; otherwise, the user is notified with the completion of all processes.
@@ IV Case Study
In this case study, we examine the proposed framework and demonstrate how the LLM is integrated into a multi-agent framework to execute the manufacturing process.
@@ IV-A Case Study Setup
A simulated manufacturing environment is equipped with three distinct machines, each specialized for milling, drilling, and threading tasks, as highlighted in Figure 3. To evaluate the system’s adaptability to new product information, we structured the case study into two distinct processes: a 2-step process and a more comprehensive 4-step process. In the 2-step process, a prototype product, referred to as ‘product-1’, is subjected to contouring and drilling operations. The 4-step process extends this sequence by incorporating counterboring and threading operations, testing the system’s capability to handle extended manufacturing workflows. The G-code specifications for these processes are adapted from [34], as detailed in Figure 4. Notably, explicit labels for G-code specifications (e.g. Step1-Contouring, Step2-Drilling) are not provided to the LLM agent. Leveraging LLM’s capability to interpret and assign appropriate G-code based on tool specifications for each resource agent demonstrates the system’s potential to efficiently manage G-code assignments, demonstrating the advanced comprehension capabilities of LLMs in a simulated manufacturing context.
Each product has an LLM-enabled PA which oversees high-level decision-making for its associated product. For instance, ‘product-1’ is instructed for a 2-step process with the following instructions as:
"Make sure to get contouring to shape the external profile of the product. Then, get a drilling operation to create precise holes. This operation can precede or follow the contouring. Proceed to Exit Buffer completed. Your product name is ’product-1’. The G-code specifications are as follows: %358 N1 N10 G90 G71 G80 G40..."
Likewise, each LLM-enabled RA is responsible for operational decisions related to its specific machine. ‘Milling1’ is tasked with both contouring and counterboring to demonstrate how LLM can adaptively assign operations to the appropriate machine. ‘Drilling1’ focuses on drilling, while ‘Threading1’ manages threading tasks. Each machine is assigned specific tool numbers that are used in G-code specifications to ensure that PAs provide only the relevant G-code instructions to each RA. For example, the ‘Milling1’ machine receives the following operational prompt:
"Perform milling operations as required, including contouring the external shape and counterboring for holes. Your machine name is Milling1. The tool number for contouring T4. The tool number for counterboring is T6."
This experiment investigates the LLM’s capability to identify and allocate correct manufacturing task and their corresponding G-code specifications. It aims to enhance the adaptability by showcasing natural language communication between PA and RA, simulating human-like interaction for seamless task execution within the manufacturing environment.

### v2

@@ I Introduction
The escalating demand for customized products and the increasing complexities of production processes are driving a transformation in the manufacturing sector. This paradigm shift has required manufacturers to quickly adjust to evolving product specifications and adeptly address unforeseen operational changes. In this context, manufacturers are required to not only maintain high-quality products and cost efficiency but also be agile and responsive to a dynamically changing environment.
@@
One significant approach for enhancing adaptability and quick response of manufacturing systems is using multi-agent control [1]. A Multi-Agent System (MAS) includes several autonomous agents for decision-making, communication, and coordination [2]. Many multi-agent architectures have been developed for manufacturing systems [3, 4]. By enabling decentralized decision-making and enhancing coordination among various agents, MAS have contributed significantly to handling complexities in the manufacturing system [5].
@@
However, there are still challenges in MAS frameworks in manufacturing. While agents in MAS can automate processes and make decisions, their ability to adapt to new formats or specifications is limited, leading to increased downtime. Thus, a challenge includes the limited ability of agents to dynamically adapt to new information including product specifications (RCh1). Another challenge lies in processing unstructured textual data. Textual data is crucial because it is more expressive than numerical data, offering vast knowledge that is currently underutilized in manufacturing, where structured data fails to include varied influencing factors [6]. Natural language processing capabilities to formalize unstructured textual data can reduce downtime and optimize processes [6]. Therefore, natural language processing is essential for interpreting and implementing textual data from human operators or interconnected systems (RCh2).
@@
The advent of Large Language Models (LLMs) presents a novel solution to these challenges. LLMs, such as OpenAI’s GPT-3 [7] and GPT-4 [8], provides opportunities to bring more flexibility and adaptability. With Machine to Machine (M2M) protocols, such as Message Queuing Telemetry Transport (MQTT) [9] and MTConnect [10] that uses predefined structures, incorporating LLMs offer benefits. LLMs can interpret, generate, and act on natural language instructions, facilitating complex decision-making processes and adaptive responses to changing manufacturing conditions [11]. The integration of natural language processing capabilities may also benefit operators and other system users who have limited experience with the specific process or system [12]. Natural language processing capabilities can allow new users to quickly adapt to managing and optimizing processes and resources [13]. The main contributions of this paper include: Leveraging LLM to allocate and execute manufacturing tasks based on user instructions to enhance task efficiency (Cn1), enabling LLM agents to adapt to evolving specifications ensuring the production process is executed with the appropriate resources (Cn2), and applying natural language processing capabilities to enable agents to comprehend context-specific instructions from users and other agents (Cn3).
@@
The rest of the manuscript is organized as follows. Section II provides background regarding MAS, LLMs in MAS, and LLMs in manufacturing. Section III details the functioning of the proposed framework. Section IV demonstrates a case study to illustrate the applications and effectiveness of our framework. Section V presents challenges and limitations while integrating LLM. Section VI summarizes the paper and discusses future work.
@@ II-A Multi-Agent Systems in Manufacturing
Various MAS architectures have been developed to enhance agent communication, significantly contributing to the increased manufacturing system flexibility. For instance, the Axiomatic Design of a Multi-Agent Reconfigurable Mechatronic System (ADMARMS) framework facilitates design-based reconfigurability [14]. Studies by Wang et al. [15] and Kim et al. [16] integrated MAS with big data analytics and reinforcement learning. Additionally, Kovalenko et al. [17] and Bi et al. [4] focused on dynamic resource and task allocation within MAS, with frameworks for resource task negotiation and a model-based resource agent architecture to address manufacturing disruptions.
While these MAS architectures were developed to enhance agent communication for increased flexibility, they focus on utilizing predefined algorithms for task allocation and agent communication. Furthermore, there has been limited exploration in using natural language interpretation from human operators or other agents to enhance the responsiveness of manufacturing process control. By leveraging LLMs, agents can improve operational flexibility and efficiency in addressing these gaps.
@@ II-B Large Language Models in Multi-Agent Systems
The emergence of LLMs has sparked interest in developing interacting autonomous agents powered by the LLM technology [18]. LLM-enabled agents demonstrate enhanced problem-solving capabilities through multi-agent discussions and diverse applications. For instance, Park et al. [19] showcase a simulation with 25 generative agents in a virtual town, facilitating social understanding studies. Another work by Wu et al. [20] introduces an open-source framework called AutoGen that facilitates customizable agent interactions and conversation programming. ChatEval [21] uses debater agents in an LLM-powered MAS for text evaluation, achieving enhanced accuracy and depth over traditional methods and single agent setups. MetaGPT [22], a meta-programming framework, enhances multi-agent collaboration using an assembly line approach for software engineering tasks.
These works demonstrate advanced multi-agent interactions, realistic behavior simulation, and task efficiency, which can enhance adaptability, coordination, and advanced natural language processing capabilities in manufacturing. Building on these insights, we explore integrating LLMs in a multi-agent framework to address the unique requirements and benefits in the manufacturing sector.
@@ II-C Large Language Models in Manufacturing
With the advancement of LLM technology, recent studies have been increasingly exploring both the opportunities and challenges of implementing LLMs within the manufacturing domain. Makatura et al. [11] explores the application of LLMs in design and manufacturing, including CAD/CAM. Their approach investigates the potential role of LLMs in a variety of tasks, including design specification and manufacturing instruction, and highlights the capabilities and limitations. The analysis of GPT-4 shows strengths in design and manufacturing knowledge but faces challenges in reasoning and accuracy, with potential solutions including specialized languages and better context management [11]. Jignasu et al. [23] provide a comprehensive evaluation of LLMs for understanding, debugging, and manipulating G-code in 3D printing, testing their capabilities in error detection, correction, and geometric transformations. Similar work by Badini et al. [24] explores using ChatGPT to optimize the G-code generation in additive manufacturing. These studies prove that LLM can interpret G-code, essential for tasks like debugging, manipulating, and comprehending the code. Fakih et al.[25] and Koziolek et al. [26] focus on enhancing programming in industrial control systems, specifically Programmable Logic Controllers (PLCs), with frameworks like LLM4PLC and studies assessing GPT-4’s capability to generate control logic from natural language prompts.
Several studies have integrated LLMs into manufacturing, including CAD/CAM, G-code, and PLCs, to enhance efficiency. However, these studies have not focused on using LLMs with MAS for autonomous task allocation, adaptability to evolving specifications, and interpretation of context-specific instructions. The framework described in Section III leverages LLM to enhance multi-agent interactions and decision-making for manufacturing processes. This framework develops autonomous agents that use context-specific natural language instructions to dynamically adjust to new product information and coordinate tasks in real-time. Our approach is an initial step toward enhancing the CAD/CAM workflow, automatically adapting CAM processes based on changes in CAD specifications, ensuring seamless manufacturing operations.
@@ III Framework
The novel architecture in Figure 1 is designed based on a MAS that includes both Product Agents (PAs) and Resource Agents (RAs). A PA is an agent responsible for decision-making related to a product, while an RA serves as a controller for a specific resource, such as a robot or machine [27].
Figure 1 provides an overview of an LLM-enabled multi-agent manufacturing system framework, where users initialize agents with instructions and capabilities. PAs are assigned specific processes including G-code specifications, while possessing the ability to communicate and access historical data for informed decision-making. RAs are responsible for the process execution of tasks, informed by instructions and tool specifications. LLM is applied for processing natural language inputs from both PAs and RAs. For evaluation, the chosen LLM is GPT-4, the latest and most advanced version of GPT developed by OpenAI [8].
@@ III-A Agent Initialization
The agent initialization process configures a MAS by defining roles, capabilities, and communication protocols for each agent. This setup ensures agents operate effectively and collaborate towards system goals. The following initial prompt is given to all agents in the system:
@@
These instructions ensure each agent operates within its capabilities, engages in effective communication, and contributes to the MAS’s objectives.
Following the initial prompts, the initialization process assigns specific roles to agents using various parameters. These parameters are selected to achieve adaptability, real-time response, autonomous decision-making, and efficient coordination within the MAS. Key parameters include:
  • Functions: Capabilities within the manufacturing that range from retrieving history to executing tasks.
  • Annotation: Agent’s role and responsibilities, which aid in selecting the appropriate LLM agent.
@@
  • Inbox: The agent receives and manages instructions, updates, and feedback from other agents.
@@
Integrating parameters into LLM-enabled MAS for manufacturing extends standardized approaches (e.g. FIPA [28] and JADE [29]) by incorporating natural language processing for more flexible agent interactions. Integrating user-defined instructions and parameters in agent initialization can be further expanded through user studies to ensure industry applicability.
@@
Based on the initialization information, the Agent Creation function, cp. Algorithm 1, dynamically generates agents. The Agent Creation function provides advanced LLM agents with different capabilities in manufacturing systems for adaptability to changing requirements. The algorithm iterates over each entry in the Agent_list, which contains initialization information. For every entry, configuration data configuration for the agent is loaded from a JSON, containing details such as the agent’s name, role, capabilities, and any specific functions. For each agent, an empty list function_list is prepared. If the ‘functions’ key is present, the algorithm iterates through each function name listed. For each function name, a reference function_reference is obtained by searching within a provided set of agent functions Agent_functions. If the function is found, its reference is appended to function_list. A new agent instance agent is created with the specified name, agent data, and the list of function references. This agent is then added to the Agents list. This algorithm facilitates the flexible creation of agents allowing for adaptation to various tasks.
@@ III-B Agent Capabilities and LLM Integration
Through the integration of LLM, agents gain the capability to interpret complex instructions and specifications, offering a flexible approach to manufacturing operations. The framework uses LLM to decode user instructions for manufacturing and leverages the OpenAI Application Programming Interface (API) feature called function calling [30]. This feature allows for the automatic execution of user-defined functions, including specific manufacturing operations (e.g., milling, drilling) through descriptions that include operation parameters and required inputs like process name and product specifications (e.g. G-code).
The ‘chat’ function, cp. Algorithm 2, facilitates interaction among agents, based on available function information and executing function calls as defined by the manufacturing process requirements. The outcome, function_response, includes the details like operation type, product, and machine names, along with a message, msgs, to facilitate the subsequent steps in the process. This approach enables context-aware execution of tasks, enhancing adaptability and efficiency.
@@ III-C Operational Workflow
The workflow, as depicted in Figure 2, shows how user instructions in textual data can be converted into actionable operations. This approach enables the translation of high-level directives of users into machine-executable actions, enhancing the responsiveness to changing conditions.
The workflow begins with the user initializing PAs and RAs. Users describe product goals, sequences of operations, and specifications. This information is then processed by the LLM, which generates responses to guide the process. The PA determines the suitable RA for the next task, prompting it to carry out the required operation. The selected RA constructs the necessary arguments to request the corresponding function call, leading to the execution of the specified process. Upon successful operation, the LLM formulates an appropriate response, informing the PA of the process status. This feedback updates the LLM-integrated PA with the current status, determining the completion of the task. If additional tasks remain, the cycle repeats; otherwise, the user is notified of the completion of all processes.
@@ IV Case Study
In this case study, we examine the proposed framework and demonstrate how the LLM is integrated into the framework to execute the manufacturing process.
@@ IV-A Case Study Setup
A simulated manufacturing environment includes three machines specialized for milling, drilling, and threading tasks, as highlighted in Figure 3. To evaluate the system’s adaptability, we structured the case study into two distinct processes: a 2-step process and a more comprehensive 4-step process. In the 2-step process, a prototype product, referred to as ‘product-1’, is subjected to contouring and drilling operations. The 4-step process adds counterboring and threading operations, testing the system’s capability to handle extended workflows. The G-code specifications are adapted from [31], as detailed in Figure 4. Notably, explicit labels for G-code specifications (e.g. Step1-Contouring, Step2-Drilling) are not provided to the agent. Leveraging LLM’s capability to interpret and assign appropriate G-code based on tool specifications for each RA demonstrates the system’s potential to manage G-code assignments, demonstrating the capabilities of LLMs.
Each product has an LLM-enabled PA which oversees high-level decision-making for its associated product. For instance, ‘product-1’ is instructed for a 2-step process with the following instructions:
"Make sure to get contouring to shape the external profile of the product. Then, get a drilling operation to create precise holes. Proceed to Exit Buffer when completed. Your product name is ’product-1’. The G-code specifications are as follows: %358 N1 N10 G90 G71 G80 G40..."
Each LLM-enabled RA handles decisions for its specific machine. ‘Milling1’ is tasked with both contouring and counterboring to demonstrate how LLM can assign operations adaptively. ‘Drilling1’ focuses on drilling, while ‘Threading1’ manages threading. Machines are assigned tool numbers used in G-code to ensure PAs provide relevant G-code to each RA. For example, the ‘Milling1’ machine receives this operational prompt:
"Perform milling operations as required, including contouring the external shape and counterboring for holes. Your machine name is Milling1. The tool number for contouring is T4. The tool number for counterboring is T6."
This experiment investigates the LLM’s capability to identify and allocate correct manufacturing task and their corresponding G-code specifications. It aims to enhance the adaptability by showcasing natural language communication between PA and RA, simulating human-like interaction for task execution.